Run Coordinator's Report for Tuesday 15 December Swing through
Wednesday 16 December Day
Summary
There are several open shifts 19-22 December:
19 December owl leader,
20 December owl leader,
21 December owl leader,
22 December owl leader,
22 December owl TO,
22 December owl worker.
The beam delivery for physics will end at 07:00 on
December 22.
Please sign up for the open shifts, if you are able!
During Tuesday swing shift, we continued production for SLUG 24.
Wednesday owl continued production for SLUG 24, until about 02:00.
At that time, beam dropped due to trouble with 1L14. Beam was recovered
and production running resumed at 07:00.
At the beginning of day shift, we were continuing production for
SLUG 24.
At 10:40, we started setting up for a Moller measurement. There
were several
problems getting the Moller DAQ systems up, and some trouble in the
machine
which delayed the start of Moller running. The day shift ended with the
Moller measurement still in progress, but almost complete.
We achieved 15.3
hours of ABU from Tuesday swing through the end of Wednesday day.
At the end of Wednesday day, SLUG 24 had 692.9k good pairs (as of
run 14709).
Run Plan
Continue taking production data on the LD2 target.
We will not change the IHWP overnight, but will remain in SLUG 24
until morning.
We will stop taking beam at 07:00 to prepare for the spectrometer move.
The shift leader should request a full rad. survey of the hall; we
will be changing the hall to Restricted access.
Before the hall status changes to Restricted access, the Work
Coordinator
must put up guards on the target windows. Ed was expecting to
check in
about 07:30, if OPS is ready to change Hall A to Restricted before then,
get in contact with Ed.
The survey group will be entering the hall as early as they can,
to begin
surveying the LHRS position.
We will move the RHRS from 20 degrees to 12.9 degrees. This
will require
crews of ~6 collaborators to operate the drill motors.
After the survey crew completes their measurements, we will also move the
LHRS to 12.9 degrees.
We should change the spectrometer momentum settings while we are
moving the
spectrometers.
The LHRS setting should be (RES kine #3):
P0=-4.0 GeV/c (Note: Left Q3 can only be set to -3.66 GeV/c, use this
value. All other magnets must set to -4.0 GeV/c).
The RHRS setting should be (RES kine #5):
P0=-3.10 GeV/c
We also should check the threshold settings for the DAQ while we
are in the hall.
There may be other DAQ tasks which will need to be done during
this period.
If you need to do work in the hall, please check in with the shift crew
on your
way into the hall.
We intend to try to close the hall by the end of day shift on
Thursday.
Before the hall closes, we should move the target to the "empty"
position.
After the hall closes, and beam is restored, we should do the long
beam recovery
procedure, but we will take optics target and aluminum background
measurements
before we start production running on LD2.
See the resonance run plan for more details.
Requests
There are several open shifts 18-22 December:
19 December owl leader,
20 December owl leader,
21 December owl leader,
22 December owl leader,
22 December owl TO,
22 December owl worker.
The beam delivery for physics will end at 07:00 on
December 22.
Please sign up for open shifts, if you are able!
Shift Leaders: please write a detailed shift summary
(
here is a great example)
and submit the BTA for each shift.
Please log CODA reboots in the halog and shift summary.
3rd person: Please make sure to start and stop the HRS DAQ at the
same time as
the parity DAQ and to do the prompt analysis and the HRS analyzer
for every run.
3rd person: please keep a careful eye on the online data display.
Most problems
can be caught quickly this way.
Please make sure that the shift checklist is done once each shift.